# MoralCLIP: Contrastive Alignment of Vision-and-Language Representations with Moral Foundations Theory

### v1

@@ 6. Conclusions and Opportunities
As AI systems permeate everyday life, the ability to understand moral dimensions becomes not just valuable but essential. In this paper, we introduced the first framework for multimodal moral interpretation, advancing our understanding of ethical content across visual and textual media. The key contributions and take away lessons are as follow:
@@
V&L alignment with MFT.    Overall, our results demonstrate that our moral training approach aligns the embedding space with morally-relevant dimensions, enabling alignment with moral content rather than superficial visual traits. While this doesn’t solve all aspects of moral understanding, it represents a meaningful shift in what the model attends to. This reorientation toward moral dimensions represents progress in developing multimodal systems capable of moral reasoning, highlighting the potential of value-aligned embedding spaces for future advances in this domain.

### v2

@@ 6. Conclusions and Opportunities
As AI systems permeate everyday life, the ability to understand moral dimensions becomes essential. In this paper, we introduced the first framework for multimodal moral interpretation, advancing our understanding of ethical content across visual and textual media. The key contributions and takeaway lessons are as follows:
@@
V&L alignment with MFT.   Overall, our results demonstrate that our moral training approach aligns representations with morally-relevant dimensions, enabling recognition of moral content rather than superficial visual traits. While this doesn’t solve all aspects of moral understanding, it represents a meaningful shift in what the model attends to. This reorientation toward moral dimensions represents progress in developing multimodal systems capable of moral reasoning, highlighting the potential of value-aligned embedding spaces for future advances in this domain. Future work could further explore moral similarity metrics beyond the Jaccard index that better capture nuanced relationships between moral foundations that discrete set overlap measures cannot fully represent.
Weak Labeling. While our automated labeling approach via the Visual Moral Compass enables scalable dataset creation, it introduces potential noise compared to expert annotation. However, our human evaluation study demonstrates that our classifier achieves reasonable agreement with human annotators across most moral foundations. Additionally, our approach leverages SMID’s expert-validated annotations as the foundation for training the Visual Moral Compass, ensuring our automated labels build upon established moral ground truth. Even so, future work would benefit from direct human moral annotation of visual and textual content.
@@
Acknowledgments
This work is supported by UID/04516/NOVA Laboratory for Computer Science and Informatics (NOVA LINCS) with the financial support of FCT/IP. It was also supported by the AMALIA project, funded by FCT/IP in the context of measure RE-C05-i08 of the Recovery and Resilience Program.